Using the method of Shokri and Beyler, calculate the radiant heat flux from a 10 ft diameter pool fire to a wall that is 15 ft away from the edge of the fire.

Calculate the time it would take for someone to feel pain if they were being subjected to 3 kW/m2.
Answer #1

Shokri and Beyler method :

Radiant heat flux q" = 15.4(L/D)-1.59( kW/m2)

Where L is the distance from the edge of the fire.

D is the diameter of the pool fire.

Step1: written all data what is given.

Diameter of pool fire D = 10 ft

Distance from the edge of the fire L = 15 ft away

Step2: put all datas into the formula by using unit consistency and convert all the units in M.K.S system.

D = 10 ft * (0.305m/1 ft) = 3.05 m

L = 15 ft*(0.305 m/1ft) = 4.575 m

Heat flux,

q" = 15.4*(4.575/3.05)-1.59 = 8.08 kW/m2

Time required for someone to feel pain,

t = n*(35/qr")1.33

n is the safety factor,

n = 1/2 , if qr"=< 6 kW/m2

n = 1/4 if qr"> 6 kW/m2

Here qr"= 3 kW/m2

Then, n = 1/2

t = 1/2 *(35/3)1.33 = 13.12 s

Ans: time required for someone to feel pain t = 13.12 s

radiative heat flux q" = 8.08 kW/m2